# Spencer Unger, Narmada Varadarajan, Felix Weilacher

### v1

@@ 3. Asymptotic Dimension
Recall that $\Gamma\curvearrowright X$ is an action with Schreier graph $G$ with respect to some generating set $S$ for $\Gamma$.
@@
The asymptotic dimension of $G$ is the least $d$ such that $G$ admits such a witness at every scale $r$, or $\infty$ if there is no such $d$. We write $d=\textnormal{asdim}(G)$. If $G$ is the Cayley graph of $\Gamma$ we write $d=\textnormal{asdim}(\Gamma)$.
@@
In the past few years, consideration of Borel asymptotic dimension witnesses has played a key role in Borel combinatorics and dynamics following its introduction by Conley et al. [4]. In fact, in many situations these witnesses can be chosen to be clopen, including for actions of $\mathbb{Z}^{d}$ on 0-dimensional spaces:
@@
Let $\mathbb{Z}^{d}\curvearrowright X$ be a continuous free action on a 0-dimensional second countable Hausdorff space $X$. Then $\textnormal{asdim}_{\mathbf{\Delta}^{0}_{1}}(X)=d$.
Theorem 10.7 actually applies to a much wider class of countable amenable groups $\Gamma$ with $d=\textnormal{asdim}(\Gamma)$. This includes groups of polynomial growth, polycyclic groups, and more. We will say $\Gamma$ admits clopen asymptotic dimension witnesses if $\Gamma$ satisfies the conclusion of this Theorem for some $d<\infty$. By Lemma 10.3 from [4] $d$ can always be taken to be $\textnormal{asdim}(\Gamma)$.
The following lemma touches on the connection between descriptive combinatorics and local algorithms, and is clear to any familiar with that connection. The general phenomenon, proved separately by Bernshteyn and Seward [2] is that for a “local problem”, having continuous solutions is equivalent to being able to produce solutions with efficient “local algorithms”. To avoid introducing the theory surrounding these ideas, our statement of the lemma and its proof are somewhat awkward.
@@
Suppose $\Gamma$ admits clopen asymptotic dimension witnesses and let $r\in\mathbb{N}$. The free part of the Bernoulli shift of $\Gamma$ has a clopen partition $U^{0},\ldots,U^{d}$ which is an asdim witness at scale $r$. Let $D\in\mathbb{N}$ be a bound on the diameters of the components of $G^{\leq r}\upharpoonright U^{i}$.
@@
Suppose $\textnormal{asdim}(\Gamma)=d$ and $\Gamma$ admits clopen asymptotic dimension witnesses. Suppose $\Gamma\curvearrowright X$ is free and continuous and $X$ is a compact Polish space with $\mathcal{B}$. For any $r\in\mathbb{N}$, there exists an asymptotic dimension $d$ witness $U^{0},\ldots,U^{d}$ with each $U^{i}\in\textnormal{Bool}_{\Gamma}(\mathcal{B})$.
@@ 4. Bounded Geometry Decompositions
Recall that $\Gamma\curvearrowright X$ is an action with Schreier graph $G$ with respect to some generating set $S$ for $\Gamma$. In this section we use witnesses to asymptotic dimension to construct the objects below.
@@
This definition is very close to the definition of “Weakly orthogonal decomposition” from [6], but we have dropped their additional “weak orthogonality” condition, hence the name change. Weakly orthogonal decompositions are an intermediate object in Gao et al.’s construction of Borel toasts for actions of $\mathbb{Z}^{d}$. In the next section we will see that BGDs can play the same role by essentially the same proof, so in this section we focus on their construction.
@@
(2): Immediate from the Claim, since $Y_{n}\subseteq Y_{n}^{\infty}$. Note that the diamter bound is the same $D_{n}$ used to choose $r_{n+1}$.
@@ 5. Construction of a toast from a BGD
Recall that $\Gamma\curvearrowright X$ is an action with Schreier graph $G$ with respect to some generating set $S$ for $\Gamma$.

### v2

@@ 3. Asymptotic Dimension
Recall that $\Gamma\curvearrowright X$ is an action with Schreier graph $G$ with respect to some finite generating set $S$ for $\Gamma$.
@@
The asymptotic dimension of $G$ is the least $d$ such that $G$ admits such a witness at every scale $r$, or $\infty$ if there is no such $d$. We write $d=\textnormal{asdim}(G)$. If $G$ is the Cayley graph of $\Gamma$ we write $d=\textnormal{asdim}(\Gamma)$. (This turns out not to depend on the choice of $S$.)
@@
In the past few years, consideration of Borel asymptotic dimension witnesses has played a key role in Borel combinatorics and dynamics following its introduction in [4]. In fact, in many situations these witnesses can be chosen to be clopen, including for actions of $\mathbb{Z}^{d}$ on 0-dimensional spaces:
@@
Let $\mathbb{Z}^{d}\curvearrowright X$ be a continuous free action on a 0-dimensional second countable Hausdorff space $X$ with Schreier graph $G$. Then $\textnormal{asdim}_{\mathbf{\Delta}^{0}_{1}}(G)=d$.
Theorem 10.7 actually applies to a much wider class of finitely generated amenable groups $\Gamma$.
Definition 3.3.
We say $\Gamma$ admits clopen asymptotic dimension witnesses if there is some $d\in\mathbb{N}$ such that for any continuous free action $\Gamma\curvearrowright X$ on a 0-dimensional second countable Hausdorff space $X$ with Schreier graph $G$, $\textnormal{asdim}_{\mathbf{\Delta}^{0}_{1}}(G)=d$.
By Lemma 10.3 from [4] this $d$ can always be taken to be $\textnormal{asdim}(\Gamma)$. For example, Theorem 3.2 says that $\mathbb{Z}^{d}$ admits clopen asymptotic dimension witnesses. In [4] groups of polynomial growth, polycyclic groups, and more are also shown to have this property.
The following lemma touches on the connection between descriptive combinatorics and local algorithms, and is clear to any familiar with that connection. The general phenomenon, proved separately by Bernshteyn and Seward [2, Theorem 1.14, 1.15] is that for a “local problem”, having continuous solutions is equivalent to being able to produce solutions with efficient “local algorithms”. To avoid introducing the theory surrounding these ideas, our statement of the lemma and its proof are somewhat awkward.
@@
Suppose $\Gamma$ admits clopen asymptotic dimension witnesses and let $r\in\mathbb{N}$. The free part of the Bernoulli shift of $\Gamma$ (with base space $2^{\omega})$ has a clopen partition $U^{0},\ldots,U^{d}$ which is an asdim witness at scale $r$. Let $D\in\mathbb{N}$ be a bound on the diameters of the components of $G^{\leq r}\upharpoonright U^{i}$.
@@
Suppose $\textnormal{asdim}(\Gamma)=d$ and $\Gamma$ admits clopen asymptotic dimension witnesses. Suppose $\Gamma\curvearrowright X$ is free and continuous and $X$ is a compact Polish space with basis $\mathcal{B}$. Then $\textnormal{asdim}_{\textnormal{Bool}_{\Gamma}(\mathcal{B})}(G)=d$.
@@ 4. Bounded Geometry Decompositions
Recall that $\Gamma\curvearrowright X$ is an action with Schreier graph $G$ with respect to some finite generating set $S$ for $\Gamma$. In this section we use witnesses to asymptotic dimension to construct the objects below.
@@
This definition is very close to the definition of “Weakly orthogonal decomposition” from [6], but we have dropped their additional “weak orthogonality” condition, hence the name change. Weakly orthogonal decompositions are an intermediate object in Gao, Jackson, Krohne, and Seward’s construction of Borel toasts for actions of $\mathbb{Z}^{d}$. In the next section we will see that BGDs can play the same role by essentially the same proof, so in this section we focus on their construction.
@@
(2): Immediate from the Claim, since $Y_{n}\subseteq Y_{n}^{\infty}$. Note that the diameter bound is the same $D_{n}$ used to choose $r_{n+1}$.
@@ 5. Construction of a toast from a BGD
Recall that $\Gamma\curvearrowright X$ is an action with Schreier graph $G$ with respect to some finite generating set $S$ for $\Gamma$.